# Optimizing Quantum Transformation Matrices: A Block Decomposition Approach for Efficient Gate Reduction

### v1

@@ Abstract
This paper introduces an algorithm designed to approximate quantum transformation matrix with a restricted number of gates by using the block decomposition technique. Addressing challenges posed by numerous gates in handling large qubit transformations, the algorithm provides a solution by optimizing gate usage while maintaining computational accuracy. Inspired by the Block Decompose algorithm, our approach processes transformation matrices in a block-wise manner, enabling users to specify the desired gate count for flexibility in resource allocation. Simulations validate the effectiveness of the algorithm in approximating transformations with significantly fewer gates, enhancing quantum computing efficiency for complex calculations. This work contributes a practical tool for efficient quantum computation, bridging the gap for scalable and effective quantum information processing applications.
@@ I Introduction
The block decomposition algorithm introduced by Ref. Yuan et al. (2020) has demonstrated remarkable success in sparse optimization. Due to its unique combinatorial approach, this method surpasses other optimization methods, such as alternating direction methods Song and Xia (2016); Xu et al. (2016) or decomposition methods Hamidi Ghalehjegh et al. (2010); Zarabie and Das (2019); Lu and Zhang (2012) on similar problems. Building upon the strengths of this method, our objective is to improve the efficiency of quantum circuits by carefully choosing the most precise gates to represent the circuit through exhaustive optimization for gate selection. By breaking down the transformation into manageable blocks, we aim to significantly reduce the number of gates required, making complex computations feasible and scalable.
In this paper, we focus on developing an algorithm that approximates a given transformation operator as a limited number of gates using the block decomposition method. The key objective is to bridge the gap between the success of the block decomposition algorithm in sparse optimization, providing a new approach to addressing the challenges with large qubit manipulation. By integrating these two areas, we aim to provide a practical solution that improves the efficiency and feasibility of quantum computations.
The remainder of this paper is organized as follows: We begin by presenting the mathematical formulations of the problem in Section II. Our block decomposition method for searching a limited number of gates to model the target transformation matrix is presented in Section III. In Section IV, we present our comparative analysis and numerical results. Finally, we conclude the paper in Section V with a summary of our work and future research directions.
@@ II Problem Formulation
where $\theta$, $\phi$, and $\lambda$ are the corresponding Euler angles. Appendix A provides insights on the positioning of these unitary matrices within the transformation matrix and their relationship with the qubits targeted by the gates.
@@
Note that $\|\mathbf{Y}-\mathbf{U}\|_{2}^{2}$ is defined as the square of the Euclidean norm of $\mathbf{Y}-\mathbf{U}$, where the $\frac{1}{2}$ factor is added for convenience in calculations Man and Wang (2024). This expression commonly serves as a fidelity term in optimization contexts. By optimizing Eq. (2), we obtain an optimized $\mathbf{Y}$ that can approximate the transformation operator, $\mathbf{U}$, formed by a series of $\mathbf{Y}_{k}^{\dagger}$.
For simplicity, we define $\mathbf{X}_{k}=\mathbf{Y}_{k}^{\dagger}$, such that $\mathbf{Y}=\mathbf{X}_{1}\mathbf{X}_{2}\mathbf{X}_{3}\dots\mathbf{X}_{M}=\prod_{k=1}^{M}\mathbf{X}_{k}$. Each $\mathbf{X}_{k}$ represents a $2^{n}\times 2^{n}$ decomposed matrix from $\mathbf{Y}$, suitable for representation as a quantum gate in a quantum circuit, which can each be converted into the corresponding $2\times 2$ sub-matrix,$\mathbf{x}_{k}$. We term these $\mathbf{X}_{k}$ matrices as gate matrices and restrict the number of $\mathbf{X}_{k}$ to a maximum of $M$. This limitation on $\mathbf{X}$ allows the transformation matrix $\mathbf{Y}$ to be constructed with a desired number of gates essential for a specific transformation, ensuring manageable complexity in quantum computation. By expanding $\mathbf{Y}$, we can rewrite Eq. (2) as:
@@ III Block decomposition method on gate minimizing
6:     Increment $t$ by 1.
7:  end while
@@ III.1 Choice of $\mathbf{X}_{w}$
Since $\mathbf{X}_{w}$ have a definite form, after choosing the working gate matrix for each iteration, we exhaustively search for optimal $\mathbf{X}_{w}$ that can obtain minimum objective value of Eq. (6). Note that after the section of $\mathbf{X}_{w}$, all other gate matrices $\mathbf{X}_{\neq w}$ are fixed for the update, considering as the constant $\mathbf{A}$ and $\mathbf{B}$ respectively in Eq. (6).
@@ III.2 Updating $\mathbf{X}_{w}$
The key of the proposed algorithm lies in minimizing the objective function, Eq. (6), for the gate matrix $\mathbf{X}_{w}$ in each iteration. This is achieved through an exhaustive search for the optimal $\mathbf{X}_{w}$ that minimizes the objective function, performed as an inner loop where we examine all possible choice of $\mathbf{X}_{w}$. Due to the specific structure of $\mathbf{X}_{w}$, this process is equivalent to solving for the sub-matrix $\mathbf{x}_{w}$ within $\mathbf{X}_{w}$. In each iteration, two pieces of information regarding $\mathbf{x}_{w}$ are updated, the position within $\mathbf{X}_{w}$ and the value of these elements.
Note that $\mathbf{D}$ represents a dictionary matrix that govern all elements in both $\mathbf{X}$ and $\mathbf{I}$ to be the same, in exception of elements at the positions of $(i,i)$,$(i,j)$,$(j,i)$ and $(j,j)$. We can choose the position of $\mathbf{x}_{w}$ based on the choice of $\mathbf{D}$. Through adjusting the $\mathbf{D}$ in the equality constraint, $\mathbf{D}\mathbf{X}=\mathbf{D}\mathbf{I}$, we can determine the choice of the four elements within $\mathbf{X}_{w}$ that do not correspond to an identity matrix, thus the position of $\mathbf{x}_{w}$.
@@
Furthermore, since $\mathbf{D}$ and $I$ is known, we can represent their multiplication term as $C$, rewriting Eq. (6) as:
@@ IV.1 Approximation performance
Number of Gates M=$5$ M=$10$ M=$15$ M= $20$ M=$25$
Converged value of $L$ $5.06$ $3.88$ $3.53$ $3.10$ $2.83$
No. of iterations required for convergence $12$ $40$ $58$ $56$ $58$
Time required for convergence $7.35$ $7.98$ $8.73$ $10.16$ $12.75$
@@
Number of Qubits $3$ $4$ $5$
Time required for each iteration $<1$ second $<6$ seconds $>30$ minutes
Table 2: Algorithm performance across different number of qubits.
This subsection aims to showcase the performance of the proposed algorithm by selecting different numbers of gates to approximate a given transformation matrix. Numerous simulations were conducted to ensure result consistency and algorithm convergence. The study focuses on a 3-qubit system, where any transformation can be depicted by a circuit with a maximum of $28$ gates Man and Wang (2024). In this simulation, we construct random target unitary matrices $\mathbf{U}$ by multiplying $28$ unitary $\mathbf{U}_{k}$ gates randomly generated in Matlab and approximate $\mathbf{U}$ with transformation matrices $\mathbf{Y}$ containing only $M$ gates through our algorithm. Table 1 compares the behavior and effectiveness of the approximation with varying values of $M$. Each data point in the tables represents an average of 30 trials.
@@
A promising solution to address this bottleneck involves the introduction of indicator variables for dictionary selection. This strategic implementation enables the transformation of quadratic programming sub-problems into binary quadratic problems with linear constraints Hu and Sotirov (2021). This modification reduces computational complexity significantly, leading to faster processing times and more efficient computation.
@@ IV.2 Circuit Formulation
In this section, we present an illustration involving a 3-qubit system, showcasing the process of approximating a random transformation matrix and constructing the corresponding quantum circuit. Beginning with a randomly generated transformation matrix in MATLAB, denoted as matrix $\mathbf{U}$ given in Eq. (19), with detail magnitude visually shown in Fig. 1, we translate this matrix into a quantum circuit depicted in Fig. 2. Each gate within this circuit is composed of a sequence of $R_{z}$, $R_{y}$, and $R_{z}$ gates, with the specific rotational angles detailed in Table 3. This example serves as a concrete illustration of the algorithm’s capability in reducing quantum gates in real-world applications.
@@
$U=\begin{bmatrix}-0.108&-0.231&-0.045&0.018&0.070&0.424&0.713&0.489\\ -0.113&-0.235&-0.523&-0.436&-0.295&-0.434&-0.119&0.423\\ 0.560&-0.016&-0.355&-0.494&0.416&0.165&0.142&-0.308\\ 0.490&0.387&0.368&-0.295&-0.492&0.186&-0.057&0.329\\ -0.116&0.438&0.078&-0.119&-0.089&-0.544&0.637&-0.253\\ -0.342&-0.273&0.599&-0.620&0.242&-0.035&-0.071&-0.027\\ -0.447&0.682&-0.251&-0.190&0.278&0.293&-0.190&0.192\\ 0.303&0.100&0.185&0.209&0.591&-0.434&-0.077&0.527\\ \end{bmatrix}$ (19)
Figure 2: General quantum circuit representation of a 3-qubit transformation matrix.
Gate $Rz$ $Ry$ $Rz$
Gate 1 $0.808$ $0.808$ $3.142$
Gate 2 $1.299$ $1.299$ $3.142$
Gate 3 $0.7$ $0.7$ $3.142$
Gate 4 $0.151$ $0.151$ 00
Gate 5 $0.418$ $0.418$ 00
Gate 6 $0.488$ $0.488$ 00
Gate 7 $0.308$ $0.308$ $3.142$
Gate 8 $1.556$ $1.556$ $3.142$
Gate 9 $0.658$ $0.658$ 00
Gate 10 $0.867$ $0.867$ 00
Gate 11 $0.24$ $0.24$ $3.142$
Gate 12 $0.863$ $0.863$ 00
Gate 13 $0.105$ $0.105$ 00
Gate 14 $0.131$ $0.131$ 00
Gate 15 $0.912$ $0.912$ 00
Gate 16 $0.877$ $0.877$ $3.142$
Gate 17 $0.489$ $0.489$ 00
Gate 18 $0.197$ $0.197$ $3.142$
Gate 19 $0.045$ $450.0$
Gate 20 $1.252$ $1.252$ $3.142$
Gate 21 $0.456$ $0.456$ $3.142$
Gate 22 $0.227$ $0.227$ 00
Gate 23 $1.086$ $1.086$ 00
Gate 24 $0.508$ $0.508$ $3.142$
Gate 25 $0.711$ $0.711$ $3.142$
Gate 26 $1.155$ $1.155$ $3.142$
Gate 27 $0.684$ $0.684$ $3.142$
Gate 28 $0.145$ $0.144$ 00
@@
Example 1: Position $i$ Position $j$ Corresponding Gate $Rz$ $Ry$ $Rz$ $1$ $4$ Gate 3 $1.485$ $0.688$ 00 $2$ $6$ Gate 11 $0.918$ $1.493$ $3.141$ $2$ $5$ Gate 10 $0.867$ $0.867$ 00 $2$ $8$ Gate 13 $0.105$ $0.105$ 00 $1$ $3$ Gate 2 $0.767$ $0.085$ $3.142$ $3$ $8$ Gate 18 $0.197$ $0.197$ $3.142$ $3$ $4$ Gate 14 $0.89$ $0.938$ 00 $2$ $5$ Gate 10 $1.29$ $0.478$ $3.142$ $5$ $7$ Gate 24 $0.824$ $1.332$ 00 $2$ $7$ Gate 12 $1.228$ $0.186$ 00
Example 2: Position $i$ Position $j$ Corresponding Gate $Rz$ $Ry$ $Rz$ $1$ $7$ Gate 6 $1.32$ $1.343$ $3.142$ $5$ $7$ Gate 24 $1.396$ $0.577$ $3.142$ $2$ $6$ Gate 11 $0.24$ $0.24$ $3.142$ $2$ $8$ Gate 13 $0.105$ $0.105$ 00 $3$ $4$ Gate 14 $0.131$ $0.131$ 00 $3$ $5$ Gate 15 $0.912$ $0.912$ 00 $3$ $8$ Gate 18 $0.197$ $0.197$ $3.142$ $2$ $4$ Gate 9 $0.65$ $0.502$ 00 $2$ $6$ Gate 11 $0.225$ $1.523$ 00 $1$ $2$ Gate 1 $0.213$ $0.178$ 00
@@
Following the sequence and angles of the gates outlined in Table 4, we can derive an approximation circuit $\mathbf{Y}$ of the circuit corresponding to matrix in Eq. (19), with details of the corresponding complete circuit shown in Table 3. Each individual gate is visually represented in Fig. 2, illustrating the structure of gates used to build a circuit for a 3-qubit system.
Note that Table 4 presents two examples of the converged result for our matrix approximation in Eq. (19), minimizing the number of gates from the circuit shown in Table 3. Rather than having a straightforward single global minimum, this problem is characterized by multiple local optima. However, our algorithm consistently reaches an optimal point for the loss function in Eq. (5). This was validated through a series of $30$ trials, ensuring robustness and reliability in the results obtained. The algorithm repeatedly converged with the loss function pinpointed at $3.773$ for the transformation matrix in Eq. (19), showcasing the consistency and precision of our approach.
The behavior of our algorithm is directly influenced by the initial gate selection for optimization, as detailed in Table 4. While both examples reach the same optimal loss function, they do so through different gate orders and values. Notably, certain gates, such as Gates 10, 13, and 18 in example 1, and Gates 11, 13, 14, 15, and 18 in example 2, remain unchanged. This phenomenon arises from their placement in the later stages of the block decomposition process. As prior gates have already been updated based on their values, these specific gates retain their original rotation settings.
@@
Applying $\mathbf{U}$ on state Eq. (20), $U|W\rangle$ is
$\begin{split}U|W\rangle=-0.120|000\rangle-0.608|001\rangle+0.026|010\rangle+0.152|011\rangle\\ +0.247|100\rangle+0.328|101\rangle+0.409|110\rangle+0.506|111\rangle\;\;.\end{split}$ (21)
@@
$\begin{split}\mathbf{Y}_{0}|W\rangle=-0.026|000\rangle+0.941|001\rangle-0.052|010\rangle-0.051|011\rangle\\ +0.061|100\rangle+0|101\rangle-0.281|110\rangle+0.165|111\rangle\end{split}$ (22)

### v2

@@ Abstract
This paper introduces an algorithm designed to approximate quantum transformation matrix with a restricted number of gates by using the block decomposition technique. Addressing challenges posed by numerous gates in handling large qubit transformations, the algorithm provides a solution by optimizing gate usage while maintaining computational accuracy. Inspired by the Block Decompose algorithm, our approach processes transformation matrices in a block-wise manner, enabling users to specify the desired gate count for flexibility in resource allocation. Simulations validate the effectiveness of the algorithm in approximating transformations with significantly fewer gates, enhancing quantum computing efficiency for complex calculations.
@@ I Introduction
The block decomposition algorithm introduced by Ref. Yuan et al. (2020) has demonstrated remarkable success in sparse optimization. Due to its unique combinatorial approach, this method surpasses other optimization methods, such as alternating direction methods Song and Xia (2016); Xu et al. (2016) or decomposition methods Hamidi Ghalehjegh et al. (2010); Zarabie and Das (2019); Lu and Zhang (2012) on similar problems. Building upon the strengths of this method, our objective is to improve the efficiency of quantum circuits by carefully choosing the most precise gates to represent the circuit through exhaustive optimization for gate selection. By breaking down the transformation into manageable blocks, we aim to significantly reduce the number of gates required to construct given transformation matrices.
In this paper, we focus on developing an algorithm that approximates a given transformation operator as a limited number of gates using the block decomposition method. The key objective is to bridge the gap between the success of the block decomposition algorithm in sparse optimization, providing a new approach to addressing the challenges with large qubit manipulation. By integrating these two areas, we aim to introduce flexibility in gate construction while maintaining the ability to limit the gate count.
The remainder of this paper is organized as follows: We begin by presenting the mathematical formulations of the problem in Section II. Our block decomposition method for searching a limited number of gates to model the target transformation matrix is presented in Section III. In Section IV, we present our comparative analysis and numerical results. Section V delves into the significance of our work. Finally, we conclude the paper in Section VI with a summary of our work and future research directions.
@@ II Problem Formulation
where $\theta$, $\phi$, and $\lambda$ are the corresponding Euler angles. Here, the rotation gates $R_{z}(\theta)$ and $R_{y}(\phi)$ are defined as:
$R_{z}(\theta)=\begin{bmatrix}e^{-i\theta/2}&0\\ 0&e^{i\theta/2}\end{bmatrix}\;\;\;\;\;\text{and}\quad R_{y}(\phi)=\begin{bmatrix}\cos(\phi/2)&-\sin(\phi/2)\\ \sin(\phi/2)&\cos(\phi/2)\end{bmatrix}$, (2)
which ensures that each gate is unitary with a determinant of 1. Consequently, the overall unitary matrix $\mathbf{U}$ is also restricted to have a determinant of 1, preserving the unitary property of the transformation. Appendix A provides insights on the positioning of these unitary matrices within the transformation matrix and their relationship with the qubits targeted by the gates.
@@
Note that $\|\mathbf{Y}-\mathbf{U}\|_{2}^{2}$ is defined as the square of the Euclidean norm of $\mathbf{Y}-\mathbf{U}$, where the $\frac{1}{2}$ factor is added for convenience in calculations Man and Wang (2024). This expression commonly serves as a fidelity term in optimization contexts. Note that Euclidean norm is given by the following equation:
$\|\mathbf{X}\|_{2}=\left(\sum_{i=1}^{m}\sum_{j=1}^{n}|x_{ij}|^{2}\right)^{\frac{1}{2}}\;\;.$ (4)
By optimizing Eq. (3), we obtain an optimized $\mathbf{Y}$ that can approximate the transformation operator, $\mathbf{U}$, formed by a series of $\mathbf{Y}_{k}^{\dagger}$.
[…]
@@
In the context of gate-based quantum computing, the repeated use of certain gates could lead to inefficiencies. To address this issue, particularly in the construction of the dictionary matrix $\mathbf{D}$, we have implemented a strategy that only considers gates that have not been used by the other $X_{(M\neq W)}$ gates in the circuit. By eliminating the positions in $\mathbf{D}$ that are already used by the other M1M-1 gates $X_{(M\neq W)}$, we ensure that the selected gates are not repeated. This adjustment significantly reduces the complexity required for our algorithm. Given that quadratic programming is the most computationally expensive step of this algorithm, similar to Yuan et al. (2020), we can now proceed with M1M-1 fewer quadratic programming steps instead of considering all gates, after discarding some gates for consideration.
@@ III Block decomposition method on gate minimizing
6:  Adaptively update penalty parameters $\lambda$ and $\mu$.
7:  Increment $t$ by 1.
8:end while
@@ III.1 Choice of $\mathbf{X}_{w}$
Since $\mathbf{X}_{w}$ have a definite form, after choosing the working gate matrix for each iteration, we exhaustively search for optimal $\mathbf{X}_{w}$ that can obtain minimum objective value of Eq. (8). Note that after the selection of $\mathbf{X}_{w}$, all other gate matrices $\mathbf{X}_{\neq w}$ are fixed for the update, considered as the constant $\mathbf{A}$ and $\mathbf{B}$ respectively in Eq. (8).
@@ III.2 Updating $\mathbf{X}_{w}$
The key of the proposed algorithm lies in minimizing the objective function, Eq. (8), for the gate matrix $\mathbf{X}_{w}$ in each iteration. This is achieved through an exhaustive search for the optimal $\mathbf{X}_{w}$ that minimizes the objective function, performed as an inner loop where we examine all possible choices of $\mathbf{X}_{w}$. Due to the specific structure of $\mathbf{X}_{w}$, this process is equivalent to solving the submatrix $\mathbf{x}_{w}$ within $\mathbf{X}_{w}$. In each iteration, two pieces of information are updated regarding $\mathbf{x}_{w}$, the position within $\mathbf{X}_{w}$ and the value of these elements.
Note that $\mathbf{D}$ represents a dictionary matrix that governs all elements in both $\mathbf{X}$ and $\mathbf{I}$ to be the same, with the exception of elements at the positions of (i,i)(i,i),(i,j)(i,j),$(j,i)$, and $(j,j)$. We can choose the position of $\mathbf{x}_{w}$ based on the choice of $\mathbf{D}$. Through adjusting the $\mathbf{D}$ in the equality constraint, $\mathbf{D}\mathbf{X}=\mathbf{D}\mathbf{I}$, we can determine the choice of the four elements within $\mathbf{X}_{w}$ that do not correspond to an identity matrix, thus the position of $\mathbf{x}_{w}$.
@@
Furthermore, since $\mathbf{D}$ and $I$ are known, we can represent their multiplication term as $C$, rewriting Eq. (8) as:
@@
III.3 Update of penalty parameters
Finally, we update the penalty parameters $\lambda$ and $\mu$ by adaptively based on the gradient of the function. Namely, in each iteration, considering the gradient norm $\|\nabla f(x_{k})\|$:
If the gradient norm $\|\nabla f(x_{k})\|$ is large, $\lambda_{k+1}$ and $\mu_{k+1}$ is decreased by $s_{1}$:
$\begin{split}\lambda_{k+1}=s_{1}\lambda_{k}\\ \mu_{k+1}=s_{1}\mu_{k}\end{split}$ (21)
Whereas, if the gradient norm $\|\nabla f(x_{k})\|$ is small, $\lambda_{k+1}$ and $\mu_{k+1}$ is increased by $s_{2}$:
$\begin{split}\lambda_{k+1}=s_{2}\lambda_{k}\\ \mu_{k+1}=s_{2}\mu_{k}\end{split}$ (22)
where $0<s_{1}<1$ in (21) and $1<s_{2}<2$ in (22) are small increments and decreases for the penalty parameters. By adjusting $\lambda$ based on the magnitude of the gradient norm at each iteration, the algorithm dynamically tunes the penalty parameters to ensure efficient convergence. Decreasing $\lambda$ and $\mu$ when the gradient norm is large allows for more significant updates in such cases, while increasing $\lambda$ and $\mu$ when the gradient norm is small helps fine-tune the optimization process for smoother convergence towards the optimal solution.
@@ IV.1 Approximation performance
Number of Gates M=55 M=$10$ M=$15$ M= 2020 M=$25$
Converged value of $L$ $4.51$ $3.87$ $3.21$ $2.31$ $1.83$
No. of iterations required for convergence $68$ $110$ $151$ $137$ $128$
Time required for convergence (seconds) $5.05$ $8.16$ $16.01$ $12.08$ $12.59$
@@
Number of Qubits 33 44 55
Time required for each iteration (minutes) <1<1 <15<15 >30>30
Table 2: Algorithm performance across different number of qubits, with $M=10$.
This subsection aims to showcase the performance of the proposed algorithm by selecting different numbers of gates to approximate a given transformation matrix. Numerous simulations were conducted to ensure result consistency and algorithm convergence. The study focuses on a 3-qubit system, where any transformation can be depicted by a circuit with a maximum of $28$ gates Li and Pelejo (2014), with each of these gates capturing distinct facets of the transformation process. In this simulation, we construct random target unitary matrices $\mathbf{U}$ by multiplying $28$ unitary $\mathbf{U}_{k}$ gates randomly generated in Matlab and approximate $\mathbf{U}$ with transformation matrices $\mathbf{Y}$ containing only $M$ gates through our algorithm. Table 1 compares the behavior and effectiveness of the approximation with varying values of $M$. Each data point in the tables represents an average of 30 trials.
@@
A promising solution to address this bottleneck involves the introduction of indicator variables for dictionary selection. This strategic implementation enables the transformation of quadratic programming sub-problems into binary quadratic problems with linear constraints. This modification significantly reduces the computational complexity, leading to faster processing times and more efficient computation.
@@ IV.2 Circuit Formulation
In this section, we present an illustration involving a 3-qubit system, showcasing the process of approximating a random transformation matrix and constructing the corresponding quantum circuit. Beginning with a randomly generated transformation matrix in MATLAB, denoted as matrix $\mathbf{U}$ given in Eq. (25), with detail magnitude visually shown in Fig. 1, we translate this matrix into a quantum circuit depicted in Fig. 2. Each gate in this circuit is composed of a sequence of $R_{z}$, $R_{y}$, and $R_{z}$ gates, with the specific rotational angles detailed in Table 3. This example serves as a concrete illustration of the algorithm’s capability in reducing quantum gates in real-world applications.
@@
$U_{\text{real}}=\begin{bmatrix}-0.113&-0.041&0.223&-0.192&-0.308&-0.321&0.051&0.509\\ -0.044&-0.556&-0.142&0.216&-0.256&0.310&0.103&0.027\\ -0.057&0.183&-0.031&-0.180&0.059&-0.423&-0.199&0.062\\ 0.141&-0.582&-0.168&-0.455&0.151&0.020&-0.256&-0.035\\ 0.102&-0.074&0.187&0.347&0.280&-0.348&0.041&-0.191\\ 0.196&-0.014&0.378&0.048&0.021&0.139&0.040&-0.178\\ 0.464&0.362&-0.365&0.280&-0.094&0.152&0.057&0.267\\ 0.319&0.022&0.334&0.056&0.154&-0.025&-0.209&-0.255\end{bmatrix}$ (23)
$U_{\text{imaginary}}=\begin{bmatrix}-0.573&0.008&-0.006&0.089&0.214&-0.208&0.122&-0.011\\ -0.087&-0.195&-0.002&0.401&-0.301&-0.197&-0.138&-0.306\\ 0.186&0.029&-0.271&-0.088&-0.538&-0.423&-0.234&-0.254\\ 0.083&0.020&-0.414&-0.181&0.310&-0.058&0.073&0.026\\ 0.108&-0.309&-0.179&0.378&0.132&-0.323&0.331&0.280\\ -0.151&0.149&-0.139&-0.233&-0.230&0.055&0.610&-0.465\\ 0.038&0.001&-0.352&0.021&0.331&-0.149&-0.031&-0.278\\ -0.428&0.153&-0.242&0.258&-0.018&0.252&-0.505&-0.004\end{bmatrix}$ (24)
$U=U_{\text{real}}\;\;+\;\;U_{\text{imaginary}}$ (25)
Figure 2: General form of quantum circuit to represent 33-qubit transformation matrix $\mathbf{U}$ in Eq. (25).
Gate $\theta$ $\phi$ $\lambda$
Gate 1 $0.166$ $0.166$ $0.663$
Gate 2 $0.318$ $0.318$ $1.274$
Gate 3 $0.257$ $0.257$ $2.611$
Gate 4 $0.226$ $0.226$ $2.326$
Gate 5 $0.358$ $0.358$ 2.487-2.487
Gate 6 $0.583$ $0.583$ $3.061$
Gate 7 $0.563$ $0.563$ 2.212-2.212
Gate 8 $0.251$ $0.251$ $1.443$
Gate 9 $0.788$ $0.788$ 3.063-3.063
Gate 10 $0.329$ $0.329$ $1.832$
Gate 11 $0.167$ $0.167$ 1.258-1.258
Gate 12 $0.405$ $0.405$ 0.111-0.111
Gate 13 $0.184$ $0.184$ 1.425-1.425
Gate 14 $1.098$ $1.098$ $1.411$
Gate 15 $0.389$ $0.389$ 2.669-2.669
Gate 16 $0.776$ $0.776$ 2.629-2.629
Gate 17 $0.656$ $0.656$ 0.932-0.932
Gate 18 $0.519$ $0.519$ 2.515-2.515
Gate 19 $1.328$ $1.328$ $1.415$
Gate 20 $0.244$ $0.244$ 1.239-1.239
Gate 21 $0.608$ $0.608$ 0.209-0.209
Gate 22 $0.374$ $0.374$ 1.358-1.358
Gate 23 $0.469$ $0.469$ 0.229-0.229
Gate 24 $0.64$ $0.64$ $1.374$
Gate 25 $0.233$ $0.233$ 3.026-3.026
Gate 26 $0.592$ $0.592$ 3.049-3.049
Gate 27 $0.398$ $0.398$ 1.668-1.668
Gate 28 $1.134$ $1.134$ 1.926-1.926
@@
Example 1: Position $i$ Position $j$ Corresponding Gate $\theta$ $\phi$ $\lambda$ 44 66 Gate 20 $0.641$ $0.326$ 0.774-0.774 22 66 Gate 11 $0.519$ $0.134$ $1.362$ 44 55 Gate 19 $0.339$ $0.955$ $2.071$ 22 77 Gate 12 $0.586$ $0.752$ 2.34-2.34 55 88 Gate 25 $0.93$ $0.987$ $3.02$ 33 44 Gate 14 $0.47$ $0.542$ $1.661$ 66 77 Gate 26 $1.034$ $1.048$ $0.384$ 33 55 Gate 15 $1.377$ $0.354$ $1.704$ 55 77 Gate 24 $1.433$ $0.904$ 2.842-2.842 11 $2$ Gate 1 $0.785$ $0.785$ $1.088$
Example 2: Position $i$ Position $j$ Corresponding Gate $\theta$ $\phi$ $\lambda$ 11 $6$ Gate 5 $0.047$ $1.433$ $1.227$ 11 $4$ Gate 3 $0.78$ $0.476$ 1.333-1.333 22 44 Gate 9 $0.257$ $0.257$ $2.611$ 22 77 Gate 12 $1.199$ $0.649$ $0.894$ 22 55 Gate 10 $0.329$ $0.329$ $1.832$ 22 88 Gate 13 $0.167$ $0.812$ $1.062$ 44 55 Gate 19 $0.973$ $0.954$ 2.208-2.208 55 66 Gate 23 $1.328$ $1.328$ $1.415$ 22 66 Gate 11 $1.136$ $1.168$ $0.377$ 33 77 Gate 17 $1.186$ $1.362$ $2.529$
@@
Table 3 and Table 4 explicitly display the angles $\theta$, $\phi$, and $\lambda$ corresponding to the gates in Eq. (1), with matrix forms as defined in Eq. (2). Following the sequence and angles of the gates outlined in Table 4, we can derive an approximation circuit $\mathbf{Y}$ of the circuit corresponding to matrix in Eq. (25), with details of the corresponding complete circuit shown in Table 3. Each individual gate is visually represented in Fig. 2, illustrating the structure of the gates used to build a circuit for a 3-qubit system. Our algorithm considers all possible positions for $\mathbf{X}_{w}$ through a combinatorial search across all the gates shown in Fig. 2. Consequently, the algorithm will require more computing time for a higher number of qubits, since more gates need to be searched.
Note that Table 4 presents two examples of the converged result for our matrix approximation in Eq. (25), minimizing the number of gates of the circuit shown in Table 3. Rather than having a straightforward single global minimum, this problem is characterized by multiple local optima. However, our algorithm consistently reaches an optimal point for the loss function in Eq. (7). This was validated through a series of $30$ trials, ensuring robustness and reliability in the results obtained. The algorithm repeatedly converged with the loss function pinpointed at $3.773$ for the transformation matrix in Eq. (25), showcasing the consistency and precision of our approach.
The behavior of our algorithm is directly influenced by the initial selection of the optimization gates, as detailed in Table 4. Although both examples reach the same optimal loss function, they do so through different gate orders and values. Notably, certain gates, such as Gates 10, 13, and 18 in example 1, and Gates 11, 13, 14, 15, and 18 in example 2, remain unchanged. This phenomenon arises from their placement in the later stages of the block decomposition process. As prior gates have already been updated based on their values, these specific gates retain their original rotation settings.
@@
We consider $\mathbf{U}$ on state Eq. (26), $U|W\rangle$ is
$\begin{split}\mathbf{U}|W\rangle=&\left(-0.019+0.0191i\right)|000\rangle+\left(-0.171-0.254i\right)|001\rangle\\ &+\left(-0.099-0.602i\right)|010\rangle+\left(-0.058-0.018i\right)|011\rangle\\ &+\left(0.293+0.163i\right)|100\rangle+\left(0.253+0.139i\right)|101\rangle\\ &+\left(-0.232-0.031i\right)|110\rangle+\left(0.161+0.441i\right)|111\rangle\end{split}$ (27)
@@
$\begin{split}\mathbf{Y}_{0}|W\rangle=&\left(-0.012+0.026i\right)|000\rangle+\left(-0.020+0.090i\right)|001\rangle\\ &+\left(-0.528+0.005i\right)|010\rangle+\left(0.488-0.188i\right)|011\rangle\\ &+\left(-0.043-0.100i\right)|100\rangle+\left(-0.136+0.345i\right)|101\rangle\\ &+\left(0.454+0.153i\right)|110\rangle+\left(-0.196+0.142i\right)|111\rangle\end{split}$ (28)